OpenStudy (anonymous):

Gregory teaches martial arts. He charges a one−time processing fee of $5.00 and the cost of the classes is shown below. Let x represent the number of classes and y represent the cost of classes. Based on this information, what will it cost to take 10 classes (including processing fee)? Number of Classes, x: 1 2 3 4 Cost of Classes, y: $15 $27 $39 $51

OpenStudy (anonymous):

The little diagram is the cost of classes not including the processing fee.

jimthompson5910 (jim_thompson5910):

So the y values in the table do NOT include the processing fee of $5?

OpenStudy (anonymous):

Correct.

jimthompson5910 (jim_thompson5910):

alright thanks

jimthompson5910 (jim_thompson5910):

Notice how as x increases by 1, y increases by 12 So the slope is 12/1 = 12 This means that m = 12 When x = 1, y = 15. Use this info in y = mx+b to find b y = mx+b 15 = 12(1) + b 15 = 12 + b 15-12 = b 3 = b b = 3 So the equation that models the cost of x classes is y = 12x+3 Now if you take 10 classes, then x = 10. Plug this in for x to find y y = 12x+3 y = 12(10)+3 y = 120+3 y = 123 So 10 classes will cost $123 (this is not counting the $5 processing fee) Add on the processing fee to get 123+5 = 128 So 10 classes will cost a total of $128

jimthompson5910 (jim_thompson5910):

Sorry made a typo, but I fixed it

OpenStudy (anonymous):

Oh my gosh, thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome